### v1

@@ 3.2 ZD setting
Here, we study zero discounting $\beta=1$ under Cobb-Douglas production $\theta<1$. The condition that $\beta\theta<1$ nonetheless allows us to evaluate the key summation as follows:
@@
As we apply the above and $n\to\infty$ into (7), we have the following:
@@
where the optimal trajectory for undiscounted contributions turns out to be geometrically convergent. We can then apply $\beta=1$ to arrive at the following zero discount verstion of the optimal trajectory for undiscounted contributions:
@@
$=$ $\begin{cases}+\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta>1\\ {\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta=1\\ -\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta<1\end{cases}$
The above result indicates that if ${A}\Theta>1$, then the population value is ever increasing and $n\to\infty$ must be the optimal solution; whereas, if ${A}\Theta<1$, then $n\to\infty$ must not be optimal and the population value must be maximized at a finite horizon $n^{*}\ll\infty$. The case where ${A}\Theta=1$ turns out to be the one that the planning horizon $n$ does not matter (beyond certain length) in maximizing the population value. Case V reflects ${A}\Theta>1$ where $n^{*}=\infty$, case VII reflects ${A}\Theta<1$ where $n^{*}\ll\infty$, and case VI reflects ${A}\Theta=1$ where $n^{*}$ is any number greater than $281$.888 The number is subject to decimals rounding. Figure 4 dipicts the population value function $\mathcal{V}[n]$, the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[n]$ and that of capital intensity $k_{t}^{*}[n]$, given the planning horizon $n$, for sample ZD settings V, VI, and VII.
The final case VIII implies ZD and AK setting ($\theta=\beta=1$), where $A\Theta=A=1+\rho>1$, under complete depreciation, by virtue of (13). Thus, if $\rho>0$ then $A>1$ in which case the population value would explode as described above, and we are left with $n^{*}\to\infty$. If in turn, $A=1+\rho=1$ as the case (VIII) goes, the infinite horizon population value is evaluated as follows:
@@
This indicates that there exists a finite optimum horizon $n^{*}\ll\infty$. Note that the population value function in this setting can be specified by zero discounting $(\beta\to 1)$ the AK model’s population value function (10), as follows:
@@
The proof for the second identity, given $A=1$, is appended to Appendix 3. The third identity is subject to L’Hôpital’s rule. By the first order condition $\frac{\partial\mathcal{V}[n]}{\partial n}=0$, the optimal planning horizon is evaluated as $n^{*}=\exp(-1+\log k_{0})-1\approx 54.182$.
Figure 4: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right). Note that the population value function of the top-left panel is increasing indefinitely.
@@ 3.3 Intergenerational inequality
It seems obvious from Figure 3 (middle row) that consumption inequality among generations is worsened as we expand the planning horizon in the presence of future discounting (except for the knife-edge case III with $\log(A\beta)=0$), while it seems to improve in Figure 4 (middle row), for cases with zero discounting. To check the validity of the conjecture, we prepare Figure 5 to illustrate how Lorenz curve moves in regard to the planning horizon (i.e., $n=200,400,600$) for AK setting cases I, III, IV, and for ZD setting cases V, VI, and VII. The Lorenz curves (in Figure 5) are based on the stream of optimal consumption for a given planning horizon, namely, $c_{t}^{*}[n]$. The level of inequality increases in AK settings as the planning horizon is stretched longer. On the other hand, the level of inequality is relatively insensitive to the planning horizon for zero discounting cases, except for case VI with $\log(A\Theta)=0$ where inequality decreases along with the length of the planning horizon.
Figure 5: Lorenz curves for $n=200$ (solid line), $n=400$ (dotted line), and $n=600$ (dashed line). Top row panels (from left to right) correspond to AK settings I, III, and IV. Second row panels (from left to right) correspond to ZD settings V, VI, and VII.
For further analysis, let us consider Gini Index, a popular measure of inequality, defined on the optimal stream of consumptions $c_{t}^{*}[n]$, as follows:
@@
Figure 6 illustrates Gini Index $\mathcal{G}$ as a function of planning horizon $n$ in various parameter settings. For both figures the the underlying parameters are fixed at $k_{0}=150$ and $A=1$. The left figure pertains to AK settings ($\theta=1$), while we raise the discount rate from $\beta=0.9$ (solid line) to $\beta=0.99$ (dashed line). As we can see from the (left) panel, AK settings have the tendency to unequalize consumptions between generations as the planning horizon expands, while inequality is mitigated by less discountig the future. In contrast, the right figure indicates that ZD settings have the tendency to equalize consumptions between generations as the planning horizon expands, while inequality is enhanced by the higher output elasticity of capital.
Figure 6: The panels are planning horizon vs Gini Index i.e., $(n,\mathcal{G}[n])$ for AK model cases with different discount rates (left), and zero discounting cases with different output elasticities of capital (right). The parameters setting of the left panel is $\theta=1$, $k_{0}=150$, $A=1$, $\beta=0.9$ (solid line), and $\beta=0.99$ (dashed line). The parameters setting of the right panel is $\beta=1$, $k_{0}=150$, $A=1$, $\theta=0.9$ (solid line), and $\theta=0.95$ (dashed line).
3.4 The initial action
Let us now focus on the initial action of the optimal consumption schedule which we specify, by virtue of (7) or (21), as follows:
@@ 3.4 The initial action
since, $\beta\theta<1$. In other words, the further we take future generations into account, the more current generation must sacrifice. In any case, the initial action is monotonously decreasing in the planning horizon, and ultimately converge to the following:
@@
By taking it’s derivative with respect to $\beta$ we obtain:
@@
That is to say, the more we respect future generations (by raising $\beta$), the more current generation must sacrefice. From another perspective, we can solve the equation $c_{0}^{*}[n]=s$ for $n$ in regard to (18), where $s$ is the consumption level of subsistence, as follows:
$n=\frac{\log\left((1-\beta\theta)A(k_{0})^{\theta}\right)-\log s}{\log(\beta\theta)}$
This $n$ is the subsistence-proof number of potential generations.
Finally, for AK models, $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$ by (19), but this value will approach zero if future generations are respected as equally as possible to the current one ($\beta\to 1$). That is to say, the ultimate respect for the future generations in AK models with an infinite planning horizon induces the current generation to sacrefice ultimately, i.e., $c_{0}^{*}[\infty]\to{s}$. Alternatively, if we relax the infinite horizon assumption, initial action for AK models with ultimate respect for the future generations be:
@@
and we are left with the AK and ZD version of subsistence-proof number of potential generations $n=Ak_{0}/s$.
@@ 4 Concluding Remarks
The extent that we take into account of the entire potential generation upon creating the objective function of the population (i.e., the dynasty) is usually determined by the compounding discounting factor. Notice that the potential number of intrinsic years that are accounted by compound discounting at a rate of $r$ is as much as $\int_{0}^{\infty}e^{-{r}x}\mathrm{d}x=1/r$. As regards to our problem, compound discounting at a rate of $r$ is equivalent to binary discounting at a planning horizon $n=1/r$. That is, for example, a generational discount rate of $1$% amounts to 100 generations. Binary discounting is flexible in a sense that the planning horizon is variable, so that infinity is not precluded from an option. Moreover, it is nondiscriminatory, ensuring that living generations are treated equally. With binary discounting, however, an economic growth model requires a finite horizon dynamic optimization approach.
One important feature of binary discounting is that it hinders future generations that earn below-critical-level utilities. If the phase boundary $\log(A\beta)$ for AK setting is negative, the infinite horizon solution trajectory of utility (less the critical level), is forever declining (as shown in the middle-right panel of Figure 3) and eventually surpasses the level of subsistence. Binary discounting avoids such repugnancy by limiting the longevity of the dynasty. On the other hand, if the phase boundary $\log({A}\Theta)$ for zero (compound) discounting is negative, a finite horizon must be optimal since an infinite horizon objective function approaches a negative infinity. Note also that the bottom utilities that lie at the center of the planning horizon (as shown in the middle-right panel of Figure 4), yield increasingly negative contributions to the population value.
As regards the distribution of consumption among generations, the rate of compound discounting plays the key role. That is, the less the rate of discounting is, the more equal the consumptions become. Moreover, as the planning horizon expands, the distribution of consumption becomes more uneven in the AK setting cases, whereas vice-versa in a zero discounting setting. Our another concern was the optimal initial consumption in light of the well-being of future generations. In any case the initial consumption is decreasing in the planning horizon, and ultimately, it will depend on the value of $\beta\theta$. The optimal initial consumption for infinite planning horizon will decrease as discount rate approaches zero, and under the AK setting, initial consumption itself approaches zero until it reaches the level of subsistence.
@@
Below we put down a production function and its dual unit cost function:
@@
where $p$, $r$, and $w$ denote prices for $Y$, $K$, and $L$, respectively. The remaining parameter $A$ is called the productivity and $B$ is called the dual cost function-based productivity. Applying Shephard’s lemma to the dual function leads to the followings.
@@
By virtue of these euqations, the marginal product of capital (MPK) can be readily evaluated as follows:
@@
where, $r/w\in(0,\infty)$ equals the marginal rate of substitution between capital and labor (MRS). By comparison, we are left with $B=A\theta^{\theta}(1-\theta)^{1-\theta}\equiv A\Theta$. Moreover,
$A=\left(\frac{\text{MPK}}{\Theta}\right)\left(\frac{r}{w}\right)^{\theta-1}$ $B=\text{MPK}\left(\frac{r}{w}\right)^{\theta-1}$
On the other hand, recall the breakdown equation of total output (2) and take the partial derivative as follows:
@@
According to the study in section 3.1, we note that condition $A\beta<1$ implies the non-increasing of undiscounted contributions (or utilities less the critical level) with respect to generations. The knife-edge case where $A\beta=1$ implies that the undiscounted contributions are constant over generations. Figure 7 (left) dipicts the set of possible values for $(\theta,A)$ that satisfies $A\beta\leq 1$. Also, according to the study in section 3.2, we note that condition $A\Theta<1$ implies the existence of a finite optimal planning horizon. Otherwise, if $A\Theta>1$, the optimal planning horizon will be inifinite. The knife-edge case where $A\Theta=1$ implies that the optimal planning horizon is indeterminate. Figure 7 (right) dipicts the set of possible values for $(\theta,A)$ that satisfies $A\Theta\leq 1$.
[…]
Figure 7: The shaded area (mariginal lines not included) corresponds to the set of parameters $(\theta,A)$ where $B=A\theta^{\theta}(1-\theta)^{1-\theta}<1$
@@
The value function of Bellman equation (4) is of the following:
@@
We show this by induction. First, evaluate the value function at $t={n}$ by (4):
@@
Since $k_{{n}+1}=0$ by (2.1), $c_{n}={A}(k_{n})^{\theta}$ must be true. Also, $\mathcal{V}_{{n}+1}=0$ must be true for efficiency. Thus, we know that the final maximization is bounded, i.e.,
@@
Because $S_{n-n}=S_{0}=1$ and since ${A}$ is a constant, the proposition holds true for $t={n}$.
Suppose the proposition holds true for $t+1$. Then,
@@
Below are the corresponding first order condition and its solution:
@@
Here, we used $(\beta\theta)S_{{n}-t-1}=\beta\theta+(\beta\theta)^{2}+\cdots+(\beta\theta)^{n-t}=S_{n-t}-1$. By plugging the above solution back into the maximand we arrive at the following result:
@@
The optimal trajectory of state $k_{t}^{*}[n]$ for Bellman equation (4) is:
$\ln k_{t}^{*}[n]=\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}$
@@
As we apply $t=0$ on the above (22) and by taking the logarithm,
@@
we know that the proposition is true for $t=1$.
Suppose that the proposition is true for $t$. Then, we know by (22) that:
@@
That is, the proposition is true for $t+1$. ∎
@@
The optimal consumption trajectory $c_{t}^{*}[n]$ for Bellman equation (4) is:
@@
Obvious from Proposition 2 and (21). ∎
@@
where we used the L’Hôpital’s rule twice. Then, we know that

### v3

@@ 3.2 ZD setting
Here, we study ZD, i.e., $\beta=1$, under Cobb–Douglas production with an ordinary output elasticity of capital $\theta<1$. The condition that $\beta\theta<1$ nonetheless allows us to evaluate the key summation as follows:
@@
We apply the above and $n\to\infty$ to (7) and obtain the following:
@@
where the optimal trajectory for undiscounted contributions is geometrically convergent. We can then apply $\beta=1$ to arrive at the following zero discount verification of the optimal trajectory for undiscounted contributions:
@@
$=$ $\begin{cases}+\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B>1\\ {\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B=1\\ -\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B<1\end{cases}$
The above result indicates that if $B>1$, then the population value is ever increasing, and $n\to\infty$ must be the optimal solution. However, if $B<1$, then $n\to\infty$ must not be optimal, and the population value must be maximized at a finite horizon $n^{*}\ll\infty$. The case where $B=1$ is the knife-edge case in which the planning horizon $n$ does not matter (beyond a certain length) in maximizing the population value. Case V corresponds to $B>1$, and $n^{*}=\infty$; case VII corresponds to $B<1$, and $n^{*}\ll\infty$; and case VI corresponds to $B=1$, and $n^{*}$ is any number greater than $281$.777 The number is subject to decimal rounding. Figure 4 depicts the population value function $\mathcal{V}[n]$, along with the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[n]$ and the capital intensity $k_{t}^{*}[n]$, given a planning horizon $n$, for sample ZD settings V, VI, and VII.
The final case VIII relates to the ZD/AK setting ($\theta=\beta=1$), where $B=A=1+\rho>1$, under complete depreciation, as described in (13). Thus, if $\rho>0$, then $A>1$, in which case the population value would increase without bound, as described above, leading to $n^{*}\to\infty$. If, in turn, $A=1+\rho=1$, as in case VIII, the infinite horizon population value is evaluated as follows:
@@
This indicates that there exists a finite optimum horizon $n^{*}\ll\infty$. Note that the population value function in this setting can be specified by applying a unitary discounting factor $(\beta\to 1)$ to the population value function (10) of AK models as follows:
@@
The proof for the second identity, given $A=1$, is appended to Appendix 3. The third identity is subject to L’Hôpital’s rule. By the first-order condition $\frac{\partial\mathcal{V}[n]}{\partial n}=0$, the optimal planning horizon is evaluated as $n^{*}=\exp(-1+\log k_{0})-1\approx 54.182$.
Figure 4: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right). Note that the population value function of the top-left panel is increasing indefinitely. The parameters corresponding to each of the cases are given in Table 1.
@@ 3.3 Intergenerational inequality
Figure 3 (middle row) shows that consumption inequality across generations increases as the planning horizon extends under future discounting (except in the knife-edge case III when $\log(A\beta)=0$), whereas it seems to decrease in Figure 4 (middle row) for cases with ZD. To validate this conjecture, Figure 5 shows how the Lorenz curve shifts with respect to the planning horizon (namely, $n=200,400,600$) for AK setting cases I, III, and IV and ZD setting cases V, VI, and VII. These Lorenz curves (of Figure 5) are based on the sequence of optimal consumption levels for a given planning horizon, i.e., $c_{t}^{*}[n]$. The inequality level increases in AK settings as the planning horizon expands. On the other hand, inequality level is relatively insensitive to the planning horizon for ZD setting cases, except for case VI, when $\log B=0$, where inequality decreases as the planning horizon increases.
Figure 5: Lorenz curves for $n=200$ (solid line), $n=400$ (dotted line), and $n=600$ (dashed line). The top row panels (from left to right) correspond to AK settings I, III, and IV. The second row panels (from left to right) correspond to ZD settings V, VI, and VII.
For further analysis, let us consider the Gini index, a popular measure of inequality, defined on the basis of the optimal stream of consumption $c_{t}^{*}[n]$, as follows:
@@
Figure 6 shows the Gini index $\mathcal{G}$ as a function of the planning horizon $n$ under various parameter settings. For both panels, the underlying parameters are fixed at $k_{0}=150$ and $A=1$. The left panel corresponds to the AK settings ($\theta=1$) as the discount rate is increased from $\beta=0.9$ (solid line) to $\beta=0.99$ (dashed line). In the AK settings, consumption tends to become more unequal across generations as the planning horizon expands. However, this effect is mitigated when future discounting decreases. In contrast, the right panel indicates that ZD settings tend to equalize consumption across generations as the planning horizon expands, whereas inequality is enhanced by the higher output elasticity of capital.
Figure 6: The panels depict the planning horizon vs Gini Index, i.e., $(n,\mathcal{G}[n])$ for AK setting cases with different discount rates (left), and ZD setting cases with different output elasticities of capital (right). The parameter settings for the left panel are $\theta=1$, $k_{0}=150$, $A=1$, $\beta=0.9$ (solid line), and $\beta=0.99$ (dashed line). The parameter settings for the right panel are $\beta=1$, $k_{0}=150$, $A=1$, $\theta=0.9$ (solid line), and $\theta=0.95$ (dashed line).
3.4 Initial action
Let us focus on the initial action of the optimal consumption schedule, which we specify as follows based on (7) or (21):
@@ 3.4 Initial action
since $\beta\theta<1$. In other words, the more future generations are treated equally, the more current generation must reduce their consumption. In any case, the initial action monotonically decreases in the planning horizon and ultimately converges to the following:
@@
By taking its derivative with respect to $\beta$, we have:
@@
In other words, the more we consider future generations (by raising $\beta$), the more the current generation must reduce their consumption. From another perspective, we can solve equation $c_{0}^{*}[n]=\nu$ for $n$ with respect to (18), where $\nu$ is the well-being subsistence, as follows:
$n=\frac{\log\left((1-\beta\theta)A(k_{0})^{\theta}\right)-\log\nu}{\log(\beta\theta)}$
This $n$ is the subsistence-proof size of potential generations.
Finally, for AK models, the optimal initial action by (19) is $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$, but this value approaches zero if future generations are given equal consideration as the current generation ($\beta\to 1$). That is, in AK models with an infinite planning horizon, giving ultimate consideration to future generations requires the current generation to ultimately reduce their consumption, i.e., $c_{0}^{*}[\infty]\to{\nu}$. Alternatively, if we relax the infinite horizon assumption, the initial action for AK models with ultimate consideration for the future generations becomes:
@@
and we are left with the AK/ZD version of subsistence-proof size of potential generations, that is, $n=Ak_{0}/\nu$.
@@ 4 Concluding Remarks
The extent to which we account for the entire potential generation upon creating the objective function of the population (i.e., the dynasty) is usually determined by the compounding discounting factor. The potential number of intrinsic years that are accounted for by compound discounting at a rate of $r$ is as high as $\int_{0}^{\infty}e^{-{r}x}\mathrm{d}x=1/r$. With respect to our problem, compound discounting at a rate of $r$ is equivalent to binary discounting (which eliminates consideration for utilities before the planning horizon while fully discounting those that lie beyond it) at a planning horizon $n=1/r$. Binary discounting is flexible in the sense that the planning horizon is variable so that infinity is not precluded as an option. Moreover, it is nondiscriminatory, ensuring that existing generations are treated in ultimate equity. With binary discounting, however, the roundaboutness of production requires a finite horizon dynamic programming approach.
One important feature of our welfare evaluation on the basis of critical-level utilitarianism is that it hinders future generations that earn below-critical-level utility. If the phase boundary $\log(A\beta)$ for the AK setting is negative, the infinite horizon solution trajectory of utility (less the critical level) is forever declining (as shown in the middle-right panel of Figure 3) and eventually surpasses the level of well-being subsistence. Binary discounting avoids such repugnancy by limiting the longevity of the dynasty. On the other hand, if the phase boundary $\log B$ for ZD is negative, a finite horizon must be optimal since an infinite horizon objective function approaches negative infinity. Note also that the bottom utilities that lie at the center of the planning horizon (as shown in the middle-right panel of Figure 4) yield increasingly negative contributions to the population value.
With respect to the distribution of consumption among generations, the rate of compound discounting plays a key role. That is, the lower the rate of discounting is, the more equal the consumption becomes. Moreover, as the planning horizon expands, the distribution of consumption becomes more uneven in an AK setting, whereas vice versa in a ZD setting. Another concern was the optimal initial consumption in light of the well-being of future generations. In any case, the optimal initial consumption decreases in the planning horizon, and ultimately, it depends on the value of $\beta\theta$. If $\beta\theta\to 1$ (that is, with the AK/ZD setting), the optimal initial consumption approaches the level of ultimate sacrifice under an infinite planning horizon, Conversely, any finite consumption of the current generation determines the (finite) planning horizon, i.e., the size of future generations.
@@
Below, we write a Cobb–Douglas production function and its dual unit cost function:
@@
where $p$, $r$, and $w$ denote prices corresponding to $Y$, $K$, and $L$, respectively. The remaining parameter $A$ is referred to as the productivity, and $B$ is the cost function-based productivity. Applying Shephard’s lemma on the dual function leads to the following.
@@
On the basis of these equations, the marginal product of capital (MPK) can be readily evaluated as follows:
@@
where $r/w\in(0,\infty)$ is the marginal rate of substitution (MRS) between capital and labor. By comparison, we are left with $B=A\theta^{\theta}(1-\theta)^{1-\theta}$. Moreover,
$A=\frac{\text{MPK}}{\theta^{\theta}(1-\theta)^{1-\theta}}\left(\frac{r}{w}\right)^{\theta-1},$ $B=\text{MPK}\left(\frac{r}{w}\right)^{\theta-1}$
On the other hand, we recall the breakdown equation of the total output (2) and take the partial derivative as follows:
@@
From the study in Section 3.1, we note that the condition $A\beta<1$ implies the nonincreasing property of undiscounted contributions (or utilities less the critical level) with respect to generations. The knife-edge case where $A\beta=1$ implies that the undiscounted contributions are constant over generations. Figure 7 (left) depicts the set of possible values for $(\theta,A)$ that satisfies $A\beta<1$. Additionally, from the study in Section 3.2, we note that the condition $B<1$ implies the existence of a finite optimal planning horizon. Otherwise, if $B>1$, the optimal planning horizon will be infinite. The knife-edge case where $B=1$ implies that the optimal planning horizon is indeterminate. Figure 7 (right) depicts the set of possible values for $(\theta,A)$ that satisfies $B<1$.
Figure 7: The shaded area (marginal lines not included) corresponds to the set of parameters $(\beta,A)$ where $A\beta<1$ (left), and $(\theta,A)$ where $B=A\theta^{\theta}(1-\theta)^{1-\theta}<1$ (right).
@@
The value function of the Bellman equation (4) is as follows:
@@
We show this by induction. First, the value function at $t={n}$ is evaluated via (4):
@@
Since $k_{{n}+1}=0$ by (2.1), $c_{n}={A}(k_{n})^{\theta}$ must be true. Additionally, $\mathcal{V}_{{n}+1}=0$ must be true for efficiency. Thus, the final maximization is bounded, i.e.,
@@
Because $S_{n-n}=S_{0}=1$ and since ${A}$ is a constant, the lemma holds true for $t={n}$.
Suppose that the lemma holds true for $t+1$. Then,
@@
Below are the corresponding first-order condition and its solution:
@@
Here, we use $(\beta\theta)S_{{n}-t-1}=\beta\theta+(\beta\theta)^{2}+\cdots+(\beta\theta)^{n-t}=S_{n-t}-1$. By plugging the above solution back into the maximand, we arrive at the following result:
@@
The optimal trajectory of state $k_{t}^{*}[n]$ for the Bellman equation (4) is as follows:
$\log k_{t}^{*}[n]=\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}$
@@
As we apply $t=0$ to the above (22) and take the logarithm,
@@
We know that the lemma is true for $t=1$.
Suppose that the lemma is true for $t$. We then know by (22) that:
@@
which indicates that the lemma is true for $t+1$. Hence, the lemma follows. ∎
@@
The optimal consumption trajectory $c_{t}^{*}[n]$ for the Bellman equation (4) is as follows:
@@
This is obvious from Lemma 2 and (21). ∎
@@
where we use L’Hôpital’s rule twice. Then, we know that
@@
Compliance with Ethical Standards
The authors declare that they have no conflicts of interest.